### v1

# Identifying the magnetic genes in fully- and partially-ordered V2Al ( Cr, Mn, Fe, Co, Ni) Heusler alloys

Abstract
Multicomponent Heusler alloys exhibit various magnetic properties arising from their diverse atomic compositions and crystal structures. Identifying the general physical principles that govern these behaviors is essential for advancing their potential in spintronic applications. In this work, we combine density functional theory with atomistic Monte Carlo simulations to investigate the magnetic ground states, finite-temperature magnetic transitions, and electronic structures of fully-ordered -, -type, and partially-ordered VAl ( Cr, Mn, Fe, Co, Ni) Heusler alloys. We introduce the concept of magnetic genes, defined as V--V triangular motifs connected by the nearest-neighbor (NN) exchange interactions . Within this framework, the magnetic ground states and transition temperatures across the VAl family can be consistently understood. The magnetic order is primarily governed by the NN interactions in the triangular genes, while the transition temperatures are additionally influenced by couplings. Furthermore, the magnetic genes are still proven to be effective in our calculations on partially-ordered V2Al alloys from to -type structures. Our results suggest that the concept of magnetic genes provides a unifying principle for understanding magnetic ordering in V-based Heusler alloys and could serve as a powerful guide for exploring magnetism and designing advanced spintronic materials in a broader class of Heusler systems.
I INTRODUCTION
Heusler alloys, long regarded as prototypical spintronic materials owing to their half-metallic nature, are capable of delivering nearly spin polarization at and even well above room temperature [1, 2, 3, 4, 5, 6]. Consequently, the discovery and design of high-performance Heusler compounds remain a central objective of both experimental and theoretical research [7]. On the experimental side, typical full-Heusler alloys such as Co2CrAl and Co2TiSn have been shown to sustain ferromagnetic order up to room temperature, with Curie temperatures of approximately 335 K and 333 K, respectively [8, 9]. Over the past two decades, theoretical efforts, particularly high-throughput first-principles calculations combined with Monte Carlo or spin-model simulations, have identified dozens of full-Heusler systems exhibiting magnetic anisotropy energies exceeding together with Curie temperatures above [10, 11]. These advances have substantially expanded the library of high-performance magnets and promising high- candidates for spintronic applications [12].
In particular, full-Heusler alloys with the general formula ( transition metals) have attracted considerable attention owing to their rich magnetic properties, shape-memory behavior, and sensitivity to chemical disorder [13, 14, 15]. For instance, Co2VAl exhibits ferromagnetic ordering with a Curie temperature close to room temperature, while its decreases under applied external pressure [16]. In contrast, Mn2VGa shows ferrimagnetic ordering with a significantly higher Curie temperature of 784 K [17]. The inverse-Heusler (-type) compound Mn2CoAl also displays robust ferromagnetism, with a of 720 K at ambient pressure [18], whereas Mn2FeSi, despite sharing the same crystal structure, has an anomalously low of only 67 K [19]. Moreover, both theoretical and experimental studies of -type full-Heusler alloys have revealed complex magnetic behavior. For example, Co2FeSi exhibits a remarkably high Curie temperature of about 1100 K [20], while Fe2VAl remains paramagnetic up to room temperature. Such diverse magnetic orders are generally attributed to variations in stoichiometry and chemical disorder [21], raising fundamental questions about the key factors that govern the magnetic properties of Heusler systems.
Recently, chemical disorder has attracted growing attention as an effective means of tuning magnetic order. Disorder can substantially modify magnetic exchange interactions, influence phase stability, and reshape the overall magnetic state of a material [22, 23, 24]. For example, studies on Ru2Mn ( Sn, Sb, Ge, Si) full-Heusler alloys have demonstrated the strong impact of chemical disorder on their magnetic behavior [22], with Co2FeSn exhibiting enhanced magnetism and pronounced size-dependent disorder effects [25]. Similarly, Smith et al. recently employed DC magnetron sputtering to grow epitaxial V2FeAl thin films [26]. Although V2FeAl shows a relatively high Curie temperature experimentally, its crystal structure appears to be a mixture of the - and -type phases, complicated further by intrinsic disorder [26]. Other works have demonstrated that increasing the degree of chemical disorder leads to a pronounced suppression of spontaneous magnetization and Curie temperature [26, 27]. From a broader physical perspective, the deliberate introduction of chemical disorder not only provides an alternative strategy for tuning magnetism but also establishes a promising platform to investigate the fundamental mechanisms underlying complex magnetic orders in Heusler alloys [24].
Although the magnetic properties of full-Heusler alloys have been extensively investigated experimentally, the underlying mechanisms remain elusive due to their complex atomic compositions, multi-sublattice character, and the presence of chemical disorder. Using first-principles approaches, previous works have systematically evaluated the magnetic exchange couplings for several Heusler alloy families employing frozen-magnon and linear-response techniques [28, 29]. For vanadium-based antiferromagnets, existing studies have primarily concentrated on the ordered -type structure of VAl ( Nb, Ta) [30]. Follow-up investigations extended this work by comparing the phase with the inverse phase across a broader range of compounds. In particular, Kuroda et al. reported that -type V2NbAl and V2TaAl favor antiferromagnetic coupling, whereas in VSi ( Ti, Zr, Hf), the structure often competes with or destabilizes the phase [30]. In addition, large-scale first-principles screenings of the inverse-Heusler chemical space, where Sc, Ti, V, Cr, Mn; Al, Si, As; and spans from Ti to Zn, have identified the necessary conditions for the emergence of semimetallic states and established three Slater–Pauling rules associated with – hybridization [31]. Despite these advances, the systematic experimental synthesis and theoretical calculations of V-based VAl compounds with the complete set of Cr, Mn, Fe, Co, and Ni are still lacking. A comprehensive microscopic understanding of the mechanisms that govern magnetism in both fully- and partially-ordered V-based full-Heusler alloys is therefore essential.
In this work, we investigate the ground-state magnetic orders and electronic structures of VAl ( Cr, Mn, Fe, Co, Ni) Heusler alloys using density functional theory. For the fully-ordered -type VAl, V2CrAl and V2MnAl exhibit ferrimagnetic order, whereas the remaining compounds adopt ferromagnetic order. The situation becomes more intricate in the inverse -type VAl. While V2CrAl displays a nearly vanishing net magnetic moment with ferromagnetic order between the two different Wyckoff sites of V atoms, the other four alloys exhibit ferrimagnetic order with antiparallel spin alignment between the two V sublattices. Despite this diversity, the magnetic orders across VAl compounds can be consistently interpreted in terms of the V--V triangular pathway, governed by the NN magnetic exchange coupling . Furthermore, finite-temperature magnetic transitions were systematically investigated through atomistic Monte Carlo simulations, which reveal that the sequence of Curie temperatures among different compositions and lattice types is primarily determined by , with additional contributions from exchange interactions. Importantly, we find that the V--V triangular pathway framework remains valid even in partially-ordered systems from to -type structures. Building on this insight, we propose the concept of “magnetic genes”, defined as the V--V triangular motif connected by NN exchange interactions . We recognize that the magnetic genes could provide a promising concept for exploring the magnetic properties in more Heusler alloys.
II METHODS
To perform the structural relaxation calculations, we use the VASP code [32, 33], which implements the projector-augmented wave (PAW) method [34] and is based on the GGA-PBE generalized gradient approximation [35]. A dense -mesh of 10×10×10 is employed for the structural relaxation calcuations. The plane-wave cutoff was chosen to be 425 eV. The convergence criteria for the total energy and ionic forces were set to and , respectively.
Additionally, we performed spin-polarized density-functional [36, 37] calculations using the Atomic-orbital Based Ab-initio Computation at USTC (ABACUS) package [38, 39], an open-source first-principles platform with advanced ground- and excited-state modules, and interfaces to Wannier90. We employed the SG15 optimized norm-conserving Vanderbilt (ONCV) pseudopotentials [40] and the same exchange correlation [35]. The Kohn-Sham wave functions were expanded using a linear combination of atomic orbitals corresponding to the SG15 ONCV (Version 2.0 standard orbitals) [39]. After convergence tests, the grid cutoff for numerical integrations was set to 100 Ry.
To explore the magnetic properties in V2Al, we employed the TB2J package [41], which is based on the use of Green’s functions. The real-space Heisenberg exchange parameters
where is the on-site exchange splitting and the spin-resolved Green function. The complete set defines a classical Heisenberg Hamiltonian
which was solved with the Vampire atomistic spin-dynamics code [42].
The simulated system of V2Al is a cubic box of edge length , generated by replicating the relaxed face-centered unit cell and containing about spins, which was evolved with a stochastic spin integrator. Temperatures from to in increments were sampled using equilibration steps. The Curie temperature was determined from the inflection point of the magnetization .
| Compound | Source | ||||
| -V2CrAl | 0.901 | -0.742 | 0.026 | 1.085 | This work |
| -V2MnAl | 1.802 | -2.018 | 0.068 | 1.653 | This work |
| -V2MnAl | 1.667 | -1.653 | -0.040 | 1.641 | Jerzy et al. [43] |
| -V2FeAl | 0.537 | 1.917 | -0.021 | 2.970 | This work |
| -V2FeAl | 0.362 | 1.889 | -0.014 | 2.599 | Jerzy et al. [43] |
| -V2CoAl | 1.238 | 1.136 | 0.024 | 3.636 | This work |
| -V2NiAl | 1.249 | 0.265 | -0.005 | 2.757 | This work |
III RESULTS AND DISCUSSION
III.1 Magnetic orders of fully-ordered - and -V2Al
The magnetic ordering in Heusler alloys is strongly influenced by both their atomic composition and crystallographic structure. In this work, we focus on two prototypical lattice types, namely the and structures. As established in previous studies, these two structural variants provide a well-defined platform for examining the role of chemical disorder without altering the overall structural topology [7, 14, 28]. As illustrated in Fig. 1(a), the -type full-Heusler structure (space group , No.225) with the generic formula VAl consists of V atoms occupying the 8 Wyckoff positions, while Al and atoms are located at the 4 and 4 sites, respectively [7]. By contrast, in the -type inverse-Heusler structure (space group , No.216), shown in Fig. 1(b), the atoms occupy the 4 positions that arise from the symmetry breaking of the 8 sites in the -type structure, while the V atoms are shifted to the 4 sites. From a broader perspective, the redistribution of atomic occupations between the 4 and 4 sites, driven by the evolution from to -type structures, highlights the importance of chemical disorder in determining the magnetic properties of these compounds. A detailed analysis of disorder effects will be presented in Sec. III.4.
| Compound | Source | |||||
| -V2CrAl | -0.497 | -1.397 | 1.774 | -0.026 | 0.145 | This work |
| -V2MnAl | -0.820 | 1.466 | 1.610 | 0.007 | 2.264 | This work |
| -V2MnAl | -0.689 | 1.440 | 1.340 | -0.082 | 2.009 | Jerzy et al. [43] |
| -V2FeAl | -0.447 | 2.043 | 1.179 | 0.002 | 2.777 | This work |
| -V2FeAl | -0.790 | 2.180 | 1.699 | -0.120 | 2.950 | Jerzy et al. [43] |
| -V2FeAl | -0.310 | 2.110 | 1.20 | 0.090 | 3.090 | Skaftouros et al. [31] |
| -V2CoAl | -0.404 | 1.801 | 0.556 | 0.043 | 1.996 | This work |
| -V2CoAl | -0.54 | 2.16 | 0.50 | -0.12 | 2.0 | Zhang et al. [44] |
| -V2NiAl | -0.526 | 1.527 | 0.107 | 0.039 | 1.148 | This work |
| -V2NiAl | -0.66 | 1.74 | 0.06 | -0.06 | 1.08 | Zhang et al. [44] |
We begin by examining the effect of atomic composition on the magnetic order in -type VAl ( Cr, Mn, Fe, Co, Ni), where the crystal structure is relatively simple, containing only three inequivalent Wyckoff sites compared to the more complex -type structure. As summarized in Table 1, compounds with Fe, Co, and Ni exhibit ferromagnetic ordering, whereas those with Cr and Mn display ferrimagnetic ordering characterized by antiparallel spin alignment between the and V sublattices. In particular, our calculations show that V2MnAl adopts a ferrimagnetic ground state, with magnetic moments of per V atom and per Mn atom. By contrast, V2FeAl exhibits ferromagnetic ordering with a total magnetic moment of per formula unit, in agreement with previous reports [43]. Notably, the total magnetic moments of the Fe to Ni compounds do not vary monotonically across the series. While the magnetic moments on the V sublattice increase progressively from Fe to Ni, those on the site decrease rapidly, leading to non-monotonic behavior in the total moment. These complex magnetic orders and the non-monotonic variation in total magnetization are primarily attributed to differences in the underlying electronic structures arising from the atomic composition. A detailed discussion of the interplay between electronic structure and magnetic order will be presented in Sec. III.5.
To investigate the influence of lattice type on magnetism, we next examine the -type inverse-Heusler VAl compounds, which possess a more complex crystal structure with four inequivalent Wyckoff sites compared to the -type structure. Interestingly, among these systems, only V2CrAl exhibits ferromagnetic coupling between the two inequivalent V sites, while the magnetic moments at the V-4 and -4 sites align antiferromagnetically, as summarized in Tables 2. This contrasts sharply with the comparatively simple magnetic ordering observed in the -type compounds. More specifically, apart from V2CrAl, all other -type VAl compounds ( = Mn, Fe, Co, Ni) display antiparallel alignment of the spin moments between the V atoms at the 4 and 4 sites. Our calculated magnetic moments for -type VAl ( = Mn, Fe, Co) are in good agreement with previous reports [31, 43, 44]. Furthermore, the half-metallic compounds V2MnAl and V2FeAl exhibit total magnetic moments of 2.26 and 2.78 /f.u., respectively, which are close to the ideal values of 2 and 3 /f.u. predicted by the generalized Slater–Pauling relation of , where is the total valence-electron count [31, 45]. Notably, as varies from Cr to Ni, the local magnetic moment of the site decreases monotonically (from 1.77 to 0.11 ), whereas the total moment evolves in a nonmonotonic manner. These results highlight the intricate and composition-dependent magnetic behaviors of inverse-Heusler VAl compounds, motivating the need for a more general understanding of their underlying mechanisms.
III.2 Microscopic magnetic mechanism in - and -V2Al
To gain deeper insight into the origin of magnetism in these systems, we investigate the magnetic exchange couplings () across different atomic compositions and lattice types. As discussed above, in the relatively simple -type VAl compounds, only those with Cr and Mn exhibit ferrimagnetic order, characterized by anti-parallel spin alignment between the and V atoms. Correspondingly, as shown in Figs. 2(a) and 2(b), the NN exchange coupling () between Mn (Cr) atoms at the 4 site and V atoms at the 8 site is antiferromagnetic type, which has the largest value among all the other magnetic exchange couplings in these materials. In contrast, for Fe, Co, and Ni, the NN coupling between the and V atoms is ferromagnetic, as illustrated in Fig. 2(c). Meanwhile, the exchange interactions between - atoms are significantly weaker and do not play a decisive role in stabilizing the magnetic order. The V-V interactions, however, exhibit more intricate behavior. In V2MnAl, both the NN and next-nearest-neighbor (NNN) V-V couplings, indicated by the gray line in Fig. 2(a), are ferromagnetic. However, the NN V-V coupling in V2CrAl, colored by the gray line, acquires a finite antiferromagnetic contribution, as shown in Fig. 2(b). Taken together, these results indicate that antiferromagnetic V-Mn(Cr)-V exchange pathways form triangular motifs that ultimately stabilize ferromagnetic ordering among the V sublattice. Overall, our exchange-coupling analysis provides a microscopic explanation for the emergence of complex magnetic orders in -type VAl compounds.
Here, we investigate the physical mechanisms underlying the magnetic ordering in the more complex -type VAl compounds. In these type compounds, the magnetic moments of V1 (4 site) and (4 site) are aligned antiparallel, indicating dominant antiferromagnetic exchange coupling. As shown in Figs. 2(d) and 2(e), the NN exchange couplings (highlighted in blue) are strongly antiferromagnetic ( Cr, Mn; couplings for Fe, Co, Ni are omitted for clarity) and represent the primary factor stabilizing ferrimagnetic ordering between the and V1 sublattices. In contrast, the NN exchange coupling between and V2 (4 site), as shown in purple lines in Figs. 2(e) and 2(f), evolves rapidly from antiferromagnetic to ferromagnetic as changes from Cr to Ni. This change of is consistent with the corresponding magnetic ordering from ferrimagnetic for Cr to ferromagnetic for Mn, Fe, Co, and Ni. Notably, although the NN V2-Mn interaction is weakly antiferromagnetic type, the NNN V2-Mn coupling is strongly ferromagnetic, suggesting that V2MnAl lies near a compositional boundary where the magnetic ordering transitions from ferrimagnetic to ferromagnetic with increasing atomic number.
From our analysis of magnetic orders and exchange couplings, a comprehensive understanding of the overall magnetic behavior in VAl across different atomic compositions and lattice types can be established. In -type VAl, the diverse magnetic orders can be rationalized in terms of triangular exchange pathways formed by V--V atoms, as illustrated in the inset of Fig. 2(a). In -type VAl, the underlying mechanism is similar, but an additional type of magnetic exchange coupling is present. Our results indicate that the V1--V2 triangular pathway remains a key determinant of magnetic order in -type compounds, as shown in the inset of Fig. 2(d). Changes in the type of the coupling substantially influence the magnetic ordering in these materials. Moreover, although the direct NN exchange coupling between - atoms ( Mn, Fe, Co, Ni) is antiferromagnetic type in -VAl, the triangular -V1- pathway ultimately stabilizes ferromagnetic ordering due to the intervening antiferromagnetic X-V1 couplings, as illustrated in the inset of Fig. 2(d). In summary, despite variations in atomic composition and lattice type, the site plays a central role in controlling the magnetic properties of V-based VAl compounds through triangular pathways formed by NN V and atoms. We therefore identify these motif as a candidate “magnetic genes”.
III.3 Magnetic transition temperature of - and -V2Al
The microscopic mechanisms of magnetic exchange couplings directly determine the macroscopic magnetic properties. Based on our calculated exchange couplings in VAl, we further simulate the magnetic transition temperatures across this series of Heusler compounds. As shown in Fig. 3(a), in -type VAl, V2MnAl exhibits the highest Curie temperature, reaching approximately 900K, whereas the Curie temperatures of the remaining compounds ( Cr, Fe, Co, Ni) are significantly lower, ranging from 100 K to 300 K. The exceptionally high Curie temperature of V2MnAl originates from the effective ferromagnetic coupling between V atoms mediated by the triangular V-Mn-V pathways, which involve strong antiferromagnetic NN couplings ( meV). This is further reinforced by the direct ferromagnetic V-V interactions. Remarkably, our Monte Carlo simulations reveal that the magnitude of the NN V– exchange coupling predominantly determines the magnetic transition temperature in -type V2Al. Specifically, as shown in Fig. 3(a), decreases from 900 K for V2MnAl to 100 K for V2FeAl, consistent with the corresponding NN exchange couplings , which vary from to 2.0 meV, as illustrated in Fig. 2.
The temperature dependence of magnetic phase transitions in -type VAl compounds differs markedly from that in -type structures. As shown in Fig.3(b), the Curie temperatures in -type compounds vary within a narrower range, from approximately 700 K to 300 K, compared to the broader variation in -type materials. This narrower range reflects the more complex network of magnetic exchange couplings in the inverse-Heusler structure. Specifically, V2CrAl exhibits the highest at about 700 K, while the Curie temperature gradually decreases as changes from Fe to Ni. The magnetic transition temperature of V2MnAl shows a slight anomaly, lying between the values for Co and Ni. Despite the apparent complexity of these trends, they can be rationalized using the proposed microscopic mechanism based on the triangular exchange pathways. As discussed above, the NN exchange coupling involving the -site dominates the magnetic transition behavior. For instance, the relatively large contributes to the higher in V2CrAl. Although is also sizable, the of V2MnAl is significantly lower due to the additional influence of the coupling. In V2CrAl, the ferromagnetic – interactions enhance , whereas in V2MnAl, the coupling is antiferromagnetic, leading to a notable reduction of . In the series Fe, Co, Ni, the gradual decrease of is primarily driven by the weakening of the NN coupling, which diminishes from to 0 meV.
In summary, the magnetic ground states and magnetic transition temperatures of VAl compounds can be rationalized within the framework of the V--V triangular exchange pathway and its associated NN magnetic exchange couplings, applicable to both - and -type structures. For simplicity, the NN exchange coupling primarily governs the magnetic ordering, whereas the combination of and predominantly determines the magnetic transition temperature . Although depends on both the exchange couplings and the local magnetic moments of the constituent atoms, no straightforward correlation between the magnitude of individual magnetic moments and is evident. Nevertheless, the observed trends in the dominant exchange couplings provide a practical framework to anticipate and potentially tune the Curie temperatures in these Heusler alloys.
| Structural ordering | |||||||
| () | () | () | () | () | () | f.u. | |
| 100% | -2.018 | ––– | ––– | 1.802 | 1.802 | 0.068 | 1.653 |
| 75%/25% | -1.863 | 2.936 | -0.872 | 1.184 | 1.432 | 0.033 | 1.472 |
| 50%/50% | -1.976 | 2.731 | -0.926 | 0.630 | 1.474 | 0.016 | 1.719 |
| 25%/75% | -1.593 | 1.582 | -0.595 | 0.111 | 1.338 | 0.028 | 1.679 |
| 100% | ––– | 1.610 | -0.820 | ––– | 1.466 | 0.007 | 2.264 |
III.4 Effect of chemical disorder on magnetism of partially-ordered V2Al
In VAl compounds, the NN magnetic exchange coupling plays a central role in determining both the magnetic ground states and the magnetic transition temperatures. This raises the question of whether this framework remains valid in the presence of chemical disorder. In this section, we investigate the evolution of magnetic order and magnetic transition temperatures under partial chemical disorder, introduced by progressively transforming the -type structure toward the -type structure, as illustrated in Fig. 1(c) and 1(d). The introduction of disorder significantly reduces crystal symmetry, complicating the analysis of magnetic interactions. For clarity, we focus on the -type V2MnAl, which exhibits the highest Curie temperature among its sister compounds. Notably, our calculations show that the magnetic ground states remain robust across different degrees of chemical disorder. As summarized in Tables 3, starting from the fully-ordered -type structure, sequential swapping of Mn atoms in the 4 sites with V atoms in the 8 sites results in only minor reductions in the magnetic moment of V2 (nominal 4 site), which then stabilizes as the structure approaches the -type configuration. Furthermore, despite atom swapping, the spin orientations at the corresponding Wyckoff sites are preserved. The magnetic moments of V atoms on both 4 and 4 sites decrease gradually from the to the -type structure, as do the magnetic moments of Mn atoms on their respective sites. Overall, these results indicate that the intricate network of magnetic exchange couplings in VAl is sufficiently robust to stabilize magnetic order even in the presence of substantial chemical disorder.
The magnetic transition temperature in partially-ordered V2MnAl gradually decreases as the structure evolves from to -type structure. As shown in Fig. 4(a), the Curie temperature decreases from 900 K in the fully-ordered structure to 420 K in the configuration. Within our framework, the NN magnetic exchange couplings and are the primary factors governing the magnetic transition temperature. In the partially-ordered structures studied here, swapping V and atoms directly affects the NN coupling , while other magnetic exchange couplings are also modified. A comparison between the calculated Curie temperatures and the corresponding values in Fig. 4 shows a clear correlation. The magnitude of on the swapped sites closely tracks the variation of . For instance, in 75% -V2MnAl, is reduced to meV, and further decreases to meV ( meV and meV for the other two relevant sites) in 25% -V2MnAl, compared to meV in the fully-ordered structure. The rapid reduction of from to 75% is reflected in the corresponding decrease of from 990 K to 650 K.
Interestingly, within the V2Al series, the trend of magnetic transition temperatures in V2FeAl is opposite to that observed in V2MnAl. As shown in Fig. 5(a), the Curie temperature of V2FeAl gradually decreases from 600 K to 120 K as the structure evolves from - to type. This observation raises the question of which factors primarily govern this anomalous sequence of magnetic transition temperatures. As discussed above, the NN magnetic exchange coupling between V1 and () plays a dominant role in determining . Our calculations indicate that, although the sequence of with increasing chemical order in V2FeAl is reversed relative to V2MnAl, the NN coupling remains a key factor. As shown in Fig. 5(b), the absolute value of decreases systematically from the - to the -type structure, consistent with the observed reduction in . These results confirm the validity of as the primary determinant of the Curie temperature in partially-ordered V2FeAl.
Our calculations on partially chemically disordered VAl indicate that the NN magnetic exchange coupling remains the dominant factor governing the magnetic transition temperature. Notably, during the transition from the to the -type structure, the 4 site is consistently occupied by V2 atoms, ensuring that the V-X-V magnetic triangular pathway is preserved. Consequently, the microscopic magnetic mechanism in chemically disordered VAl is still primarily governed by the V-X-V triangular pathway and the associated NN magnetic exchange couplings .
III.5 Electronic structures of fully-ordered V2Al
Based on the calculated magnetic ground states of -type VAl in Sec. III.1, the observed complex magnetic orders and non-monotonic variation of total magnetic moments with atomic composition are proposed related to their electronic structures. In -type VAl, the relatively small total magnetic moments of V2CrAl and V2MnAl originate from the anti-parallel alignment of spin moments between the V and sublattices. This behavior is clearly reflected in the calculated density of states (DOS). As shown in Fig. 6(a), the total DOS of -type V2Al exhibits a high value at the Fermi level, indicating itinerant magnetism in these compounds [46, 47, 48]. Notably, several compounds display nearly half-metallic character. As shown in Fig. 6(b), Cr and Mn atoms exhibit pronounced occupied states with sharp peaks at approximately -1 eV and -2 eV in the spin-down channel, while their spin-up channel states are minimally occupied. In contrast, for Fe, Co, and Ni, the PDOS of atoms shows significant occupation in the spin-up channel and comparatively low values in the spin-down channel. These electronic features are consistent with the observed transition in magnetic order from ferrimagnetism ( Cr, Mn) to ferromagnetism ( Fe, Co, Ni).
Furthermore, from Fe to Ni, the PDOS of atoms shifts to lower energies in both spin-up and spin-down channels, as shown in Fig. 6(b). This trend is attributed to the increasing number of valence electrons with rising atomic number. The pronounced difference between the spin-up and spin-down PDOS of Fe atoms indicates a large magnetic moment on Fe in V2FeAl, whereas the PDOS of Ni atoms in V2NiAl is nearly symmetric between spin channels, reflecting a significantly reduced magnetic moment. Examination of the V-atom PDOS from Fe to Ni reveals an increase in spin-up states below the Fermi level for V atoms in Co- and Ni-containing compounds, particularly in the energy range from -1 to 0 eV, as shown in Fig. 6(c), while the spin-down PDOS decreases in the same energy range. This behavior correlates with the observed increase in magnetic moments of V atoms from Fe to Ni. In contrast, the PDOS of Al atoms remains largely unchanged across the series in Fig. 6(d), consistent with the negligible magnetic moment on Al atoms. Taken together, the change in magnetic order from Mn to Fe, combined with the opposing trends in magnetic moments of V and atoms from Fe to Ni, accounts for the non-monotonic variation in the total magnetic moment across the V2Al series.
IV CONCLUSION
In this work, we employed density functional theory calculations combined with Monte Carlo simulations based on Heisenberg magnetic exchange couplings to investigate V2Al alloys and introduced the concept of “magnetic genes”. These magnetic genes govern both the magnetic ground states and the magnetic transition temperatures in fully- and partially-ordered compounds. The underlying physical mechanism arises from the modulation of NN magnetic exchange couplings within the V--V triangular pathways, which is influenced by atomic composition, crystal structure, and the degree of chemical disorder. Using the framework of magnetic genes, the magnetic orders can be rationalized in terms of the V--V triangular pathways formed by NN couplings. Furthermore, the magnetic transition temperatures are predominantly determined by both NN and additional magnetic exchange interactions, irrespective of whether the alloy is fully or partially ordered. We anticipate that the identification of magnetic genes in other Heusler alloys may provide a systematic and predictive approach to understanding and tuning their diverse magnetic properties.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grants No. 52371174 and No. 12204033), the Fundamental Research Funds for the Central Universities (Grant No. FRF-TP-22-097A1), the State Key Lab of Advanced Metals and Materials (Grant No. 2022Z-13), the Young Elite Scientist Sponsorship Program by BAST (Grant No. BYESS2023301) and the High-performance Computing Platform of University of Science and Technology Beijing.References
- Gupta et al. [2023] S. Gupta, S. Chakraborty, V. Bhasin, S. Pakhira, S. Dan, C. Barreteau, J.-C. Crivello, S. N. Jha, M. Avdeev, J.-M. Greneche, D. Bhattacharyya, E. Alleno, and C. Mazumdar, High spin polarization in the disordered quaternary heusler alloy femnvga, Phys. Rev. B 108, 045137 (2023).
- Venkateswara et al. [2023] Y. Venkateswara, J. Nag, S. S. Samatham, A. K. Patel, P. D. Babu, M. R. Varma, J. Nayak, K. G. Suresh, and A. Alam, FeRhCrSi: Spin semimetal with spin-valve behavior at room temperature, Phys. Rev. B 107, L100401 (2023).
- Nishihaya et al. [2025] S. Nishihaya, M. J. A. Jardine, H. S. Inbar, A. Goswami, J. T. Dong, A. N. Engel, Y.-H. Chang, C. P. Dempsey, M. Hashimoto, D. Lu, N. Marom, and C. J. Palmstrøm, Near-half-metallic state in the half-Heusler PtMnSb film on a III-V substrate, Phys. Rev. Mater. 9, 044403 (2025).
- de Groot et al. [1983] R. A. de Groot, F. M. Mueller, P. G. v. Engen, and K. H. J. Buschow, New class of materials: Half-metallic ferromagnets, Phys. Rev. Lett. 50, 2024 (1983).
- Žutić et al. [2004] I. Žutić, J. Fabian, and S. Das Sarma, Spintronics: Fundamentals and applications, Rev. Mod. Phys. 76, 323 (2004).
- Katsnelson et al. [2008] M. I. Katsnelson, V. Y. Irkhin, L. Chioncel, A. I. Lichtenstein, and R. A. de Groot, Half-metallic ferromagnets: From band structure to many-body effects, Rev. Mod. Phys. 80, 315 (2008).
- Graf et al. [2011] T. Graf, C. Felser, and S. S. Parkin, Simple rules for the understanding of Heusler compounds, Progress in Solid State Chemistry 39, 1 (2011).
- Husmann and Singh [2006] A. Husmann and L. J. Singh, Temperature dependence of the anomalous hall conductivity in the Heusler alloy , Phys. Rev. B 73, 172417 (2006).
- Shigeta et al. [2018] I. Shigeta, Y. Fujimoto, R. Ooka, Y. Nishisako, M. Tsujikawa, R. Y. Umetsu, A. Nomura, K. Yubuta, Y. Miura, T. Kanomata, M. Shirai, J. Gouchi, Y. Uwatoko, and M. Hiroi, Pressure effect on the magnetic properties of the half-metallic Heusler alloy , Phys. Rev. B 97, 104414 (2018).
- Kübler et al. [2007] J. Kübler, G. H. Fecher, and C. Felser, Understanding the trend in the curie temperatures of -based Heusler compounds: Ab initio calculations, Phys. Rev. B 76, 024414 (2007).
- Fortunato et al. [2024] N. M. Fortunato, X. Li, S. Schönecker, R. Xie, A. Taubel, F. Scheibel, I. Opahle, O. Gutfleisch, and H. Zhang, High-throughput screening of all--metal Heusler alloys for magnetocaloric applications, Chemistry of Materials 36, 6765 (2024).
- Marathe and Herper [2023] M. Marathe and H. C. Herper, Exploration of all- Heusler alloys for permanent magnets: An ab initio based high-throughput study, Phys. Rev. B 107, 174402 (2023).
- Guo et al. [2025] Y. Guo, C. Guo, Y. Xu, B. Zheng, X. Ni, and F. Tian, Machine learning investigation of chemical disorder effects in 24-valence-electron full-Heusler alloys with single magnetic element, Journal of Magnetism and Magnetic Materials 629, 173322 (2025).
- Picozzi et al. [2004] S. Picozzi, A. Continenza, and A. J. Freeman, Role of structural defects on the half-metallic character of and Heusler alloys, Phys. Rev. B 69, 094423 (2004).
- Devi et al. [2018] P. Devi, S. Singh, B. Dutta, K. Manna, S. W. D’Souza, Y. Ikeda, E. Suard, V. Petricek, P. Simon, P. Werner, S. Chadhov, S. S. P. Parkin, C. Felser, and D. Pandey, Adaptive modulation in the magnetic shape-memory Heusler alloy, Phys. Rev. B 97, 224102 (2018).
- Kanomata et al. [2010] T. Kanomata, Y. Chieda, K. Endo, H. Okada, M. Nagasako, K. Kobayashi, R. Kainuma, R. Y. Umetsu, H. Takahashi, Y. Furutani, H. Nishihara, K. Abe, Y. Miura, and M. Shirai, Magnetic properties of the half-metallic Heusler alloys and under pressure, Phys. Rev. B 82, 144415 (2010).
- Suto et al. [2024] H. Suto, V. Barwal, K. Simalaotao, Z. Li, K. Masuda, T. Sasaki, Y. Miura, and Y. Sakuraba, Negative spin polarization of Mn2VGa Heusler alloy thin films studied in current-perpendicular-to-plane giant magnetoresistance devices, Journal of Applied Physics 135, 203901 (2024).
- Ouardi et al. [2013] S. Ouardi, G. H. Fecher, C. Felser, and J. Kübler, Realization of spin gapless semiconductors: The Heusler compound , Phys. Rev. Lett. 110, 100401 (2013).
- Životský et al. [2022] O. Životský, K. Skotnicová, T. Čegan, J. Juřica, L. Gembalová, F. Zažímal, and I. Szurman, Structural and magnetic properties of inverse-Heusler Mn2FeSi alloy powder prepared by ball milling, Materials 15, 697 (2022).
- Wurmehl et al. [2005] S. Wurmehl, G. H. Fecher, H. C. Kandpal, V. Ksenofontov, C. Felser, H.-J. Lin, and J. Morais, Geometric, electronic, and magnetic structure of : Curie temperature and magnetic moment measurements and calculations, Phys. Rev. B 72, 184434 (2005).
- Singh and Mazin [1998] D. J. Singh and I. I. Mazin, Electronic structure, local moments, and transport in , Phys. Rev. B 57, 14352 (1998).
- Khmelevskyi et al. [2015] S. Khmelevskyi, E. Simon, and L. Szunyogh, Antiferromagnetism in full Heusler alloys: Effects of magnetic frustration and chemical disorder, Phys. Rev. B 91, 094432 (2015).
- Sokolovskiy et al. [2012] V. V. Sokolovskiy, V. D. Buchelnikov, M. A. Zagrebin, P. Entel, S. Sahoo, and M. Ogura, First-principles investigation of chemical and structural disorder in magnetic Ni2Mn1+xSn1-x Heusler alloys, Phys. Rev. B 86, 134418 (2012).
- Decolvenaere et al. [2019] E. Decolvenaere, E. Levin, R. Seshadri, and A. Van der Ven, Modeling magnetic evolution and exchange hardening in disordered magnets: The example of Heusler alloys, Phys. Rev. Mater. 3, 104411 (2019).
- Li et al. [2013] T. Li, J. Duan, C. Yang, and X. Kou, Synthesis, microstructure and magnetic properties of Heusler Co2FeSn nanoparticles, Micro & Nano Letters 8, 143 (2013).
- Smith et al. [2024] R. Smith, Z. Gercsi, R. Zhang, K. Siewierska, K. Rode, and J. Coey, Effects of disorder on the magnetic properties of the Heusler alloy V2FeAl, Acta Materialia 267, 119733 (2024).
- Miroshkina et al. [2022] O. N. Miroshkina, B. Eggert, J. Lill, B. Beckmann, D. Koch, M. Y. Hu, T. Lojewski, S. Rauls, F. Scheibel, A. Taubel, M. Šob, K. Ollefs, O. Gutfleisch, H. Wende, M. E. Gruner, and M. Friák, Impact of magnetic and antisite disorder on the vibrational densities of states in Heusler alloys, Phys. Rev. B 106, 214302 (2022).
- Galanakis et al. [2002] I. Galanakis, P. H. Dederichs, and N. Papanikolaou, Slater-Pauling behavior and origin of the half-metallicity of the full-Heusler alloys, Phys. Rev. B 66, 174429 (2002).
- Chico et al. [2016] J. Chico, S. Keshavarz, Y. Kvashnin, M. Pereiro, I. Di Marco, C. Etz, O. Eriksson, A. Bergman, and L. Bergqvist, First-principles studies of the Gilbert damping and exchange interactions for half-metallic Heuslers alloys, Phys. Rev. B 93, 214439 (2016).
- Kuroda et al. [2020] F. Kuroda, T. Fukushima, and T. Oguchi, First-principles study of magnetism and phase stabilities of V2 based antiferromagnetic Heusler alloys, Journal of Applied Physics 127, 193904 (2020).
- Skaftouros et al. [2013] S. Skaftouros, K. Özdoğan, E. Şaşıoğlu, and I. Galanakis, Generalized Slater-Pauling rule for the inverse Heusler compounds, Phys. Rev. B 87, 024420 (2013).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
- Kresse and Joubert [1999] G. Kresse and D. Joubert, From ultrasoft pseudopotentials to the projector augmented-wave method, Phys. Rev. B 59, 1758 (1999).
- Blöchl [1994] P. E. Blöchl, Projector augmented-wave method, Phys. Rev. B 50, 17953 (1994).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
- Hohenberg and Kohn [1964] P. Hohenberg and W. Kohn, Inhomogeneous electron gas, Phys. Rev. 136, B864 (1964).
- Kohn and Sham [1965] W. Kohn and L. J. Sham, Self-consistent equations including exchange and correlation effects, Phys. Rev. 140, A1133 (1965).
- Li et al. [2016] P. Li, X. Liu, M. Chen, P. Lin, X. Ren, L. Lin, C. Yang, and L. He, Large-scale ab initio simulations based on systematically improvable atomic basis, Computational Materials Science 112, 503 (2016), computational Materials Science in China.
- Zhou et al. [2025] W. Zhou, D. Zheng, Q. Liu, D. Lu, Y. Liu, P. Lin, Y. Huang, X. Peng, J. J. Bao, C. Cai, Z. Jin, J. Wu, H. Zhang, G. Jin, Y. Ji, Z. Shen, X. Liu, L. Sun, Y. Cao, M. Sun, J. Liu, T. Chen, R. Liu, Y. Li, H. Han, X. Liang, T. Bao, N. Chen, H. Ren, X. Zhang, Z. Liu, Y. Fu, M. Liu, Z. Li, T. Wen, Z. Tang, Y. Xu, W. Duan, X. Wang, Q. Gu, F.-Z. Dai, Q. Zheng, J. Zhao, Y. Zhang, Q. Ou, H. Jiang, S. Liu, B. Xu, S. Xu, X. Ren, L. He, L. Zhang, and M. Chen, ABACUS: An electronic structure analysis package for the AI era, arXiv e-prints , 2501.08697 (2025).
- Schlipf and Gygi [2015] M. Schlipf and F. Gygi, Optimization algorithm for the generation of ONCV pseudopotentials, Computer Physics Communications 196, 36 (2015).
- He et al. [2021] X. He, N. Helbig, M. J. Verstraete, and E. Bousquet, TB2J: A python package for computing magnetic interaction parameters, Computer Physics Communications 264, 107938 (2021).
- Skubic et al. [2008] B. Skubic, J. Hellsvik, L. Nordström, and O. Eriksson, A method for atomistic spin dynamics simulations: implementation and examples, Journal of Physics: Condensed Matter 20, 315203 (2008).
- Goraus et al. [2023] J. Goraus, M. Fijałkowski, J. Czerniewski, and W. Gumulak, Magnetic properties of V2MnGa, V2MnAl, V2FeGa and V2FeAl, Journal of Magnetism and Magnetic Materials 585, 171103 (2023).
- Zhang et al. [2012] X. Zhang, X. Dai, G. Chen, H. Liu, H. Luo, Y. Li, W. Wang, G. Wu, and G. Liu, First-principle study for full-Heusler compounds V2YAl (Y=V, Cr, Mn, Fe, Co, Ni) and a discussion to Slater–Pauling rule, Computational Materials Science 59, 1 (2012).
- Faleev et al. [2017] S. V. Faleev, Y. Ferrante, J. Jeong, M. G. Samant, B. Jones, and S. S. P. Parkin, Unified explanation of chemical ordering, the Slater-Pauling rule, and half-metallicity in full Heusler compounds, Phys. Rev. B 95, 045140 (2017).
- Xu et al. [2024] Y. Xu, Y.-C. Wang, X. Jin, H. Liu, Y. Liu, H. Song, and F. Tian, Mechanism of magnetic phase transition in correlated magnetic metal: insight into itinerant ferromagnet Fe3‑δGeTe2, Commun. Phys. 7, 381 (2024).
- Xu et al. [2025a] Y. Xu, X. Jin, J. Xiang, H. Zhang, and F. Tian, investigation of the magnetic phase transition in the itinerant ferromagnet , Phys. Rev. B 111, 155142 (2025a).
- Xu et al. [2025b] Y. Xu, X. Jin, H. Tang, and F. Tian, Electronic structures and magnetism in van der waals flat-band material , Europhysics Letters 151, 36002 (2025b).